# Gordian split links in the Gehring ropelength problem

### v1

@@ 5.2 Constructing an isotopically Gordian split link
Next we exhibit a thickly embedded split link $L$ fulfilling the conditions of Proposition 5.28, which has an initial length of less than $68\pi+$ 6. This means that $L$ cannot be split through a thick homotopy. A fortiori, $L$ cannot be split through a thick isotopy. So $L$ is a Gordian split link both as a thick and as a thickly embedded link. We construct the link $L=(M,A^{1},A^{2},C)$ such that the main curve $M$ is in the $XY$-plane, the auxiliary curves $A^{1}$ and $A^{2}$ are in the $XZ$-plane, and the center curve is in the $YZ$-plane. The following plots, as well as numerical calculations of some of the curves’ lengths, were created and conducted with Mathematica.
@@
There is a unique straight line segment in the quadrant $y,z\geq 0$ connecting a point on the unit circle around $G$ and a point on the boundary curve $\partial\mathcal{G}$ while being tangent to both the unit circle and $\mathcal{G}$. We call the start of this line segment $N$ and the end of it $O$. Let $P=\left(0,0,\tfrac{\sqrt{7}}{2}\right)$ and $M=\left(0,\tfrac{\sqrt{15}}{2},0\right)$. We define the first quarter of our center curve $C$ as the concatenation of the circle segment from $M$ to $N$, the straight line segment from $N$ to $O$, and the segment of the boundary curve $\partial\mathcal{G}$ from $O$ to $P$. The other three quarters of $C$ are defined by reflecting the first segment from $M$ to $P$ along both the $Y$-axis and the $Z$-axis. Mathematica was not able to solve for the coordinates of $N$ and $O$ exactly, so they were calculated numerically instead. The angle $\beta=\angle MGN$ was calculated as $\beta\approx 1.2554$. The length of the straight line segment $s$ from $N$ to $O$ was calculated as $\mathrm{len}(s)\approx 1.0925$. The length of the segment $b$ from $O$ to $P$ along $\partial\mathcal{G}$ was numerically calculated as $\mathrm{len}(b)\approx 0.2116$. The length of $C$ is therefore $\mathrm{len}(C)=4\left(\beta+\mathrm{len}(s)+\mathrm{len}(b)\right)\approx 10.2380$. Finally, we get
@@
We have constructed $L$ to be a thick link, but have not yet shown that $L$ is thickly embedded. The open radius-$\tfrac{1}{2}$ normal disk bundles around $A^{1}$, $A^{2}$ and $M$ are obviously embedded. To see that the open radius-$\tfrac{1}{2}$ normal disk bundle around $C$ is embedded, we need to check that the segment $b$ is $\mathcal{C}^{1,2}$. This is true. In fact, all of $\mathcal{G}$ is $\mathcal{C}^{1,2}$. We omit a proof as it is cumbersome and not very interesting. Numerical calculations show that the curvature of $\partial\mathcal{G}$ peaks at the points $P$ and $V$ where it is around $1.51$.
@@
The difference between the lower bound we derived in section 5.1 and the length of our constructed configuration $L$ is less than $0.08\%$ of $\mathrm{len}(L)$. While it is unlikely that $L$ can be shortened much, it is highly likely that the lower bound of $68\pi+$ 6 from Proposition 5.28 can be significantly improved. A careful examination of the proof of Proposition 5.28 shows that we never needed to assume that the distances between $A^{1}$ and $C$ and between $A^{2}$ and $C$ stay at least $1$ throughout the homotopy. In fact, we have not even used that $A^{1}$ and $A^{2}$ stay disjoint from $C$. We only needed that $(A^{1},A^{2},M)$ and $(M,C)$ stay thick links throughout the homotopy of $L$.
@@ 6.1 Circle decompositions - the idea
Figure 7: A conjoined circle decomposition: The image of the blue curve is the union of three coplanar, pairwise tangent unit circles. The image of the red curve is the union of three pairwise tangent unit circles, each piercing the centers of $2$ blue circles.
@@
In the original configuration, there are two strands going through each circle’s center. After a small perturbation, for every circle, pick one point each on both of the strands through that circle minimizing their respective distance. This gives a set of two points for every circle. These sets are marked $A,B,C,D,E,F$ in Figure 8. We now split up the two link-components into $6$ curves each, along the points making up the sets $A,B,C,D,E,F$, and pair them up as indicated by the color scheme. Let us calculate the sum $S_{\text{blue}}$ of the lengths of the blue segments. We define the closed curve $C_{\text{blue}}$ by connecting the two blue components by a straight line segment through the $2$ points in $B$ and by a straight line segment through the $2$ points in $F$. Then the two strands piercing the blue circle will pierce the center-of-mass spanning disk of $C_{\text{blue}}$ in (at least) one point each, say in the points $p$ and $q$, where $\{p,q\}=A$. We have
$\mathrm{len}(C_{\text{blue}})\geq 2\pi+2||p-q||=2\pi+2\mathrm{diam}(A),$
@@
Doing the same thing for all other colors and summing the lengths we get that all the terms involving diameters of sets $A,B,C,D,E,F$ cancel, so that the only thing remaining is
@@
This would then show that any thick link sufficiently close to the original configuration is at least as long, so the original configuration seems to be a local minimum. Besides the many missing details, there are two problems with this proof. First, even if our arguments are correct, we only get that the initial configuration is a local minimum, which is not yet sufficient to conclude that it is a sink. Secondly, we made a mistake. We do not know that the straight line segment keeps distance at least $1$ to the points $p$ and $q$ intersecting the spanning disk of $C_{\text{blue}}$. But this was a condition implicitly used when we claimed $C_{\text{blue}}\geq 2\pi+2||p-q||$. We will see that we can make adjustments to the argument to get
@@
where near 00 we have that $\varphi(x)\sim x$ to first order, but with $\varphi(x)>x$ for $x>0$. This adjusted argument is no longer strong enough to show that the original configuration was a local minimum. In some ways, the reason why the argument fails is that there are too many ”breaking points”. We can fix this by augmenting $L$.
@@
In the configuration shown in Figure 9, we go through every one of the original $6$ circles one additional time. This only introduces one new breaking point for each circle, which corresponds to one new term of $-\varphi(\mathrm{diam}(\cdot))$ for each of the sets $A,B,C,D,E,F$. At the same time, each of the circles introduces a term of $2\mathrm{diam}(\cdot)$ for the sets $A,B,C,D,E,F$. This gives the inequality
@@
with equality if and only if $\mathrm{diam}(A)=\mathrm{diam}(B)=\dots=\mathrm{diam}(F)=0$. i.e. we have $\mathrm{len}(L^{\prime})\geq 24\pi$ with equality if and only if $L^{\prime}$ is also a circle decomposition.
@@ 6.2 Circle decompositions - the details
6.36 Definition.
Let $C$ be a closed curve such that the definition domain of $C$ decomposes into $I_{1}\cup I_{2}\cup\dots\cup I_{n}$, where each $I_{i}$ is a compact interval of positive length with $\max(I_{i})=\min(I_{i+1})$, for all $i$ (here indices are understood to be cyclic and the definition domain is seen as an interval with endpoints identified). We define $a_{i}$ via $[a_{i},a_{i+1}]=I_{i}$. Let $\{A_{j}\}_{j=1}^{m}$ and $\{\alpha_{l}\}_{l=1}^{k}$ each be a partition of $\{1,2,\dots,n\}$ into non-empty sets. We say that $\left(C,\{a_{i}\}_{i=1}^{n},\{A_{j}\}_{j=1}^{m},\{\alpha_{l}\}_{l=1}^{k}\right)$ is a circle decomposition of $C$ if for all $i_{1},i_{2}\in\{1,\dots,n\}$ and $j\in\{1,\dots,m\}$:
  • 1.
@@
Let $(C,\{a_{i}\},\{A_{j}\},\{\alpha_{l}\})$ be a circle decomposition. If $a_{i}$ is on the boundary of one of the intervals $I_{i^{\prime}}$ with $i^{\prime}\in A_{j}$, then we say that $a_{i}$ is along the circle $C^{j}$. We say that $\alpha_{l}$ is along a circle $C^{j}$ if there is some $i\in\alpha_{l}$ with $a_{i}$ along $C^{j}$.
@@
We say that $C$ intersects $D$ transversely at a point $p=D(x)=C(y)$ if there is an open disk $U\subseteq[0,1]^{n}$ around $x$ and an open interval $I\subseteq[0,1]$ around $y$ such that $D|_{U}$ and $C|_{I}$ only intersect in $p$ and such that the Brouwer degree of $(D-C)|_{U\times I}$ around 00 is ±$\pm 1$.
@@
  • 2.
    $K_{j}$ circle of $C_{1}$implies there is some $\beta_{l}$ such that $z_{j}=C_{2}(\beta_{l}),$ and every intersection of $C_{2}$ and $D_{j}$ is transverse,
  • 3.
    $K_{j}$ circle of $C_{2}$implies there is some $\alpha_{l}$ such that $z_{j}=C_{1}(\alpha_{l}),$ and every intersection of $C_{1}$ and $D_{j}$ is transverse.
@@
  • 2.
    for every $A$-index $l$: #$\#\{j\text{ }A\text{-index }\mid l\text{ along }C_{1}^{j}\}<2\#(h^{A})^{-1}(l)$,
  • 3.
    for every $B$-index $l$: #$\#\{j\text{ }B\text{-index }\mid l\text{ along }C_{2}^{j}\}<2\#(h^{B})^{-1}(l)$
  • 4.
@@
The main tool in proving Theorem 6.44 will be finding a lower bound for the length of ”almost closed” curves avoiding some set of points. In the following lemmas, $U_{r}(\cdot)$ will refer to the open $r$-neighborhood of a set or point and $B_{r}(\cdot)$ will refer to the closed $r$-neighborhood of a set or point.
@@
Let $D_{1}$ and $D_{2}$ be two open disks in $\mathbb{R}^{2}$, with radii $r_{1},r_{2}$ and centers $z_{1},z_{2}$ respectively, with $r_{1},r_{2}\in[\tfrac{1}{2},1]$ and $||z_{1}-z_{2}||<\tfrac{1}{2}$. For all $a,b\in\mathbb{R}^{2}\setminus\left\{D_{1}\cup D_{2}\right\}$ with $||a-b||<\tfrac{1}{2}$ and all $m\in D_{1}\cap D_{2}$ not collinear with $a$ and $b$, there exists a curve $C$ from $a$ to $b$, such that
$\mathrm{len}(C)\leq\arcsin(2||a-b||)-||a-b||,$
obtained from radial projection of the straight-line segment $\overline{ab}$ from $m$.
@@
Take any $a,b\in\mathbb{R}^{2}\setminus(D_{1}\cup D_{2})$ and consider the straight line segment $\overline{ab}$. If $\overline{ab}$ avoids $D_{1}\cup D_{2}$, then we take $C=\overline{ab}$. If $\overline{ab}$ intersects $D_{1}\cup D_{2}$ then we define $C$ by radially projecting (from $m$) the part of $\overline{ab}$ that lies between the first and the final intersection of $\overline{ab}$ and $D_{1}\cup D_{2}$ onto $\partial(D_{1}\cup D_{2})$ and by leaving the rest of $\overline{ab}$ the same. Then $C=l_{1}\frown A\frown l_{2}$ where $l_{1},l_{2}$ are straight line segments which are part of $\overline{ab}$ and where $A$ is the path on the boundary. There are three cases, either $A$ is a circular arc, the concatenation of two circular arcs $A_{1}\frown A_{2}$, or the concatenation of three circular arcs $A_{1}\frown A_{2}\frown A_{3}$. Assume $A$ is a circular arc, say from $v_{1}$ to $v_{2}$ and assume without loss of generality that the arc is on $\partial D_{1}$. Then
$\mathrm{len}(C)$ $=\mathrm{len}(l_{1})+\mathrm{len}(A)+\mathrm{len}(l_{2})$
$\leq\mathrm{len}(l_{1})+\mathrm{len}(l_{2})+2r_{1}\arcsin\left(\frac{||v_{1}-v_{2}||}{2r_{1}}\right)$
$\leq 2r_{1}\arcsin\left(\frac{\mathrm{len}(l_{1})+\mathrm{len}(l_{2})}{2r_{1}}\right)+2r_{1}\arcsin\left(\frac{||v_{1}-v_{2}||}{2r_{1}}\right)$
$\leq 2r_{1}\arcsin\left(\frac{\mathrm{len}(l_{1})+||v_{1}-v_{2}||+\mathrm{len}(l_{2})}{2r_{1}}\right)$
$\leq\arcsin\left(\mathrm{len}(l_{1})+||v_{1}-v_{2}||+\mathrm{len}(l_{2})\right)$
$=\arcsin(||a-b||)$
$\leq\arcsin(2||a-b||)-||a-b||,$
where we used that for $x\geq 0$, the map $r\mapsto 2r\arcsin(\tfrac{x}{2r})$ is monotonically decreasing and that $\tfrac{1}{2}\leq r_{1}$. If $A=A_{1}\frown A_{2}$ is a union of two circular arcs, then $\partial D_{1}$ and $\partial D_{2}$ intersect in exactly two points, say $p$ and $q$. Assume without loss of generality that $\{p\}=A_{1}\cap A_{2}$ and that $\overline{ab}$ intersects $D_{1}$ before $D_{2}$. We know that $\overline{ab}$ must intersect $\partial D_{1}$ twice, say at points $w_{1},w_{3}$, and it must intersect $\partial D_{2}$ twice, say at points $w_{2},w_{4}$. These intersections happen in order $w_{1},w_{2},w_{3},w_{4}$. Then $A_{1}$ a circular arc on $D_{1}$ from $w_{1}$ to $p$ and $A_{2}$ a circular arc on $D_{2}$ from $p$ to $w_{4}$. Let $E_{1}$ be the circular arc from $p$ to $w_{3}$ on $\partial D_{1}$ and let $E_{2}$ be the circular arc from $w_{2}$ to $p$ on $\partial D_{2}$. Then $A_{1}\frown E_{1}$ is the circular arc from $w_{1}$ to $w_{3}$, hence $\mathrm{len}(A_{1}\frown E_{1})=2r_{1}\arcsin\left(\tfrac{||w_{1}-w_{3}||}{2r_{1}}\right)\leq\arcsin(||w_{1}-w_{3}||)$ and similarly $\mathrm{len}(E_{2}\frown A_{2})=2r_{2}\arcsin\left(\tfrac{||w_{2}-w_{4}||}{2r_{2}}\right)\leq\arcsin(||w_{2}-w_{4}||)$. Note that $E_{2}\frown E_{1}$ is a path from $w_{2}$ to $w_{3}$, so $||w_{2}-w_{3}||\leq\mathrm{len}(E_{2})+\mathrm{len}(E_{1})$. This yields
$\mathrm{len}(A)\leq$ $\arcsin(||w_{1}-w_{3}||)+\arcsin(||w_{2}-w_{4}||)-||w_{2}-w_{3}||$
$=$ $\arcsin(||w_{1}-w_{2}||+2||w_{2}-w_{3}||+||w_{2}-w_{4}||)-||w_{2}-w_{3}||$
$\leq$ $\arcsin(||w_{1}-w_{2}||+2||w_{2}-w_{3}||+||w_{3}-w_{4}||)$
$+\arcsin(||w_{1}-w_{2}||+||w_{3}-w_{4}||)$
$-||w_{1}-w_{2}||-||w_{2}-w_{3}||-||w_{3}-w_{4}||$
$\leq$ $\arcsin(2||w_{1}-w_{4}||)-||w_{1}-w_{4}||.$
It follows that $\mathrm{len}(C)\leq\arcsin(2||a-b||)-||a-b||$. Finally, assume that $A=A_{1}\frown A_{2}\frown A_{3}$ is the union of three circular arcs. Without loss of generality assume $\{p\}=A_{1}\cap A_{2}$, $\{q\}=A_{2}\cap A_{3}$, that $A_{1}$ and $A_{3}$ are arcs on $\partial D_{1}$ and that $A_{2}$ is an arc on $\partial D_{2}$. Then $A_{2}$ is the arc from $p$ to $q$ on $\partial D_{2}$. Let $v_{1}$ denote the start of $A_{1}$ and let $v_{2}$ denote the end of $A_{3}$ and let $E_{2}$ be the arc from $p$ to $q$ on $\partial D_{1}$. Then $A_{1}\frown E_{2}\frown A_{3}$ is an arc from $v_{1}$ to $v_{2}$ on $\partial D_{1}$. Therefore $\mathrm{len}(A_{1}\frown E_{2}\frown A_{3})=2r_{1}\arcsin\left(\tfrac{||v_{1}-v_{2}||}{2r_{1}}\right)\leq\arcsin(||v_{1}-v_{2}||)$. This yields
$\mathrm{len}(A)=$ $\mathrm{len}(A_{1}\frown E_{2}\frown A_{3})+(\mathrm{len}(A_{2})-\mathrm{len}(E_{2}))$
$\leq$ $\arcsin(||v_{1}-v_{2}||)+\left(\mathrm{len}(A_{2})-\mathrm{len}(E_{2})\right)$
$=$ $\arcsin(||v_{1}-v_{2}||)+2r_{2}\arcsin\left(\frac{||p-q||}{2r_{2}}\right)-2r_{1}\arcsin\left(\frac{||p-q||}{2r_{1}}\right)$
$\leq$ $\arcsin(||v_{1}-v_{2}||)+\arcsin(||p-q||)-||p-q||$
$\leq$ $\arcsin(||v_{1}-v_{2}||)+\arcsin(||v_{1}-v_{2}||)-||v_{1}-v_{2}||$
$\leq$ $\arcsin(2||v_{1}-v_{2}||)-||v_{1}-v_{2}||$
where we used that $x\mapsto\arcsin(x)-x$ is monotonically increasing and that $||p-q||\leq||v_{1}-v_{2}||$. This then yields $\mathrm{len}(C)\leq\arcsin(2||a-b||)-||a-b||$. ∎
@@
Let $x,y,m\in\mathbb{R}^{3}$ with $||x-y||,||x-m||,||y-m||<\tfrac{1}{2}$ Let $a,b\in\mathbb{R}^{3}\setminus U_{1}(\{x,y\})$ with $||a-b||<\tfrac{1}{2}$. Let $P$ be a plane through $a$, $b$ and $m$. Then there is a curve $C\subseteq P$ from $a$ to $b$, avoiding $U_{1}(\{x,y\})$ with $\mathrm{len}(C)\leq\arcsin(2||a-b||)-||a-b||$ and such that $C$ is obtained by radial projection of the straight line segment $\overline{ab}$ away from $m$.
@@
Toward a contradiction, assume $m\in\overline{ab}$. Then $||a-x||\leq||a-m||+||m-x||\leq||a-b||+||m-x||<1$, which is a contradiction to $||a-b||,||m-x||<\tfrac{1}{2}$. So $m\notin\overline{ab}$. This means that if $a$, $b$ and $m$ are collinear, then either $a\in\overline{mb}$ or $b\in\overline{am}$. Without loss of generality assume that $a\in\overline{mb}$. Toward a contradiction, assume there is some $c\in\overline{ab}$ with $c\in U_{1}(x)$. Since $U_{1}(x)$ is convex and since $a\in\overline{mc}$, this implies $a\in U_{1}(x)$, which is not the case. We conclude that $\overline{ab}$ and $U_{1}(x)$ are disjoint. Analogously we can conclude that $\overline{ab}$ and $U_{1}(y)$ are disjoint. We define $C=\overline{ab}$. If $a$, $b$ and $m$ are not collinear, then by taking the intersection of $U_{1}(x)$ and $U_{1}(y)$ with $P$, the problem reduces to Lemma 6.45. ∎
@@
Let $C:[0,1]\to\mathbb{R}^{3}$ be a curve. Let $\bar{C}=C\frown S$ be the closed curve obtained by concatenating $C$ with the straight line segment from $a=C(1)$ to $b=C(0)$ with $||a-b||<\tfrac{1}{2}$. Let $m$ be a point in the convex hull of $\bar{C}$, $K$ the $m$-cone of $\bar{C}$ and $x,y\in K$ with $1\leq d(x,C),d(y,C)$ and $||x-y||,||x-m||,||y-m||<\tfrac{1}{2}.$ Let $\varphi(t)=\arcsin(2t)-t$. Then
$\mathrm{len}(C)\geq 2\pi+2||x-y||-\varphi(||a-b||)$
@@
Let $P$ be a plane containing $m$, $a$ and $b$. Then, by Lemma 6.46, there is a curve $\tilde{C}$ in $P$ connecting $a$ to $b$ obtained by projecting $s$ radially away from $m$ in $P$ with $\mathrm{len}(\tilde{C})\leq F(||a-b||)$ and with $1\leq d(x,\tilde{C}),d(x,\tilde{C})$. Let $\tilde{K}$ be the $m$-cone of $C\frown\tilde{C}$. Note that $m\in\mathrm{conv}(C\frown S)\subseteq\mathrm{conv}(C\frown\tilde{C})$ and that $x,y\in\tilde{K}$. Since $C\frown\tilde{C}$ avoids $U_{1}(\{x,y\})$, we can apply Lemma 4.24 to get
$\mathrm{len}(C)+\mathrm{len}(\tilde{C})=\mathrm{len}(C\frown\tilde{C})\geq 2\pi+2||x-y||,$
@@
Let $C_{i}:[0,1]\to\mathbb{R}^{3}$, $i=1,\dots,k$ be curves. Let $C$ be the curve obtained by interpolating between curves $C_{i}$ and $C_{i+1}$ with straight line segments $s_{i,i+1}$ (here $k+1$ is interpreted as $1$). Let $m$ be in the convex hull of $C$, $K$ the $m$-cone of $C$, and let $x,y\in K$ with $1\leq d(x,C_{i}),d(y,C_{i})$ for $i=1,\dots,k$ and $||x-y||,||x-m||,||y-m||<\tfrac{1}{2}$. Let $\varphi(t)=\arcsin(2t)-t$. Then
$\sum_{i=1}^{k}\mathrm{len}(C_{i})\geq 2\pi+2||x-y||-\sum_{i=1}^{k}\varphi\left(||C_{i}(1)-C_{i+1}(0)||\right).$
The following theorem is an ”almost-closed” analogon of Proposition 4.25 and an easy consequence of the previous proposition.
@@
$\sum_{i=1}^{k}\mathrm{len}(C_{i})\geq 2\pi+2\mathrm{diam}(A)-\sum_{i=1}^{k}\varphi\left(||C_{i}(1)-C_{i+1}(0)||\right),.$
@@
$||D-\tilde{D}||,||C-\tilde{C}||<\varepsilon\implies\tilde{C}\cap\tilde{D}\neq\emptyset.$
@@
Let $C_{1},C_{2},\dots,C_{n}$ be $n$ curves, all defined on compact intervals such that $p=C_{1}(x_{1})=C_{2}(x_{2})=\dots=C_{n}(x_{n})$ is the only intersection point of $C_{1}$ and $C_{2}$. Further assume that every $C_{i}$ is an embedding. For every $\varepsilon>0$ there is a $\delta>0$ such that the following holds: Let $||C_{i}-\tilde{C}_{i}||<\delta$ for all $i$. Let $y_{1},\dots,y_{n}$ be such that
@@
$||D-\tilde{D}||,||C-\tilde{C}||<\delta,\hskip 5.69046pt\tilde{q}\in\tilde{C}\cap\tilde{D}\implies||\tilde{q}-p||<\varepsilon.$

### v2

@@ 6.2 Constructing an isotopically Gordian split link
Next we exhibit a thickly embedded split link $L$ fulfilling the conditions of Proposition 6.1, which has an initial length of less than $8\pi+6$. This means that $L$ cannot be split through a thick homotopy. A fortiori, $L$ cannot be split through a thick isotopy. So $L$ is a Gordian split link both as a thick and as a thickly embedded link. We construct the link $L=(M,A^{1},A^{2},C)$ such that the main curve $M$ is in the $XY$-plane, the auxiliary curves $A^{1}$ and $A^{2}$ are in the $XZ$-plane, and the center curve is in the $YZ$-plane. The following plots, as well as numerical calculations of some of the curves’ lengths, were created and conducted with Mathematica.
@@
There is a unique straight line segment in the quadrant $y,z\geq 0$ connecting a point on the unit circle around $G$ and a point on the boundary curve $\partial\mathcal{G}$ while being tangent to both the unit circle and $\partial\mathcal{G}$. We call the start of this line segment $N$ and the end of it $O$. Let $P=\left(0,0,\tfrac{\sqrt{7}}{2}\right)$ and $M=\left(0,\tfrac{\sqrt{15}}{2},0\right)$. We define the first quarter of our center curve $C$ as the concatenation of the circle segment from $M$ to $N$, the straight line segment from $N$ to $O$, and the segment of the boundary curve $\partial\mathcal{G}$ from $O$ to $P$. The other three quarters of $C$ are defined by reflecting the first segment from $M$ to $P$ along both the $Y$-axis and the $Z$-axis. Mathematica was not able to solve for the coordinates of $N$ and $O$ exactly, so they were calculated numerically instead. The angle $\beta=\angle MGN$ was calculated as $\beta\approx 1.2554$. The length of the straight line segment $s$ from $N$ to $O$ was calculated as $\mathrm{len}(s)\approx 1.0925$. The length of the segment $b$ from $O$ to $P$ along $\partial\mathcal{G}$ was numerically calculated as $\mathrm{len}(b)\approx 0.2116$. The length of $C$ is therefore $\mathrm{len}(C)=4\left(\beta+\mathrm{len}(s)+\mathrm{len}(b)\right)\approx 10.2380$. Finally, we get
@@
We have constructed $L$ to be a thick link, but have not yet shown that $L$ is thickly embedded. The open radius-$\tfrac{1}{2}$ normal disk bundles around $A^{1}$, $A^{2}$ and $M$ are obviously embedded. To see that the open radius-$\tfrac{1}{2}$ normal disk bundle around $C$ is embedded, we need to check that the segment $b$ is $\mathcal{C}^{1,2}$. This is true. In fact, all of $\partial\mathcal{G}$ is $\mathcal{C}^{1,2}$. We omit a proof as it is not very interesting.
@@
The difference between the lower bound we derived in Subsection 6.1 and the length of our constructed configuration $L$ is less than $0.08\%$ of $\mathrm{len}(L)$. While it is unlikely that $L$ can be shortened much, it is highly likely that the lower bound of $8\pi+6$ from Proposition 6.1 can be significantly improved. A careful examination of the proof of Proposition 6.1 shows that we never needed to assume that the distances between $A^{1}$ and $C$ and between $A^{2}$ and $C$ stay at least $1$ throughout the homotopy. In fact, we have not even used that $A^{1}$ and $A^{2}$ stay disjoint from $C$. We only needed that $(A^{1},A^{2},M)$ and $(M,C)$ stay thick links throughout the homotopy of $L$.
@@ 7.1 Circle decompositions - the idea
Figure 7: A conjoined circle decomposition: The image of the blue curve is the union of three coplanar, pairwise tangent unit circles. The image of the red curve is the union of three pairwise tangent unit circles, each piercing the centers of two blue circles.
@@
In the original configuration, there are two strands going through each circle’s center. After a small perturbation, for every circle, pick one point each on both of the strands through that circle minimizing their respective distance. This gives a set of two points for every circle. These sets are marked $A$, $B$, $C$, $D$, $E$, $F$ in Figure 8. We now split up the two link-components into $6$ curves each, along the points making up the sets $A$, $B$, $C$, $D$, $E$, $F$, and pair them up as indicated by the color scheme. Let us calculate the sum $S_{\text{blue}}$ of the lengths of the blue segments. We define the closed curve $C_{\text{blue}}$ by connecting the two blue components by a straight line segment through the $2$ points in $B$ and by a straight line segment through the $2$ points in $F$. Then the two strands piercing the blue circle will pierce the center-of-mass spanning disk of $C_{\text{blue}}$ in (at least) one point each, say in the points $p$ and $q$, where $\{p,q\}=A$. We hope to have
[…]
@@
Doing the same thing for all other colors and summing the lengths we would then get that all the terms involving diameters of sets $A,B,C,D,E,F$ cancel, so that the only thing remaining is
@@
[…]
@@
where near 0 we have that $\varphi(x)\sim x$ to first order, but with $\varphi(x)>x$ for $x>0$. This adjusted argument is no longer strong enough to show that the original configuration was a local minimum. In some ways, the reason why the argument fails is that there are too many breaking points. We can fix this by augmenting $L$.
@@
In the configuration shown in Figure 9, we go through every one of the original $6$ circles one additional time. This only introduces one new breaking point for each circle, which corresponds to one new term of $\varphi(\mathrm{diam}(\cdot))$ for each of the sets $A,B,C,D,E,F$. At the same time, each of the circles introduces a term of $2\mathrm{diam}(\cdot)$ for the sets $A,B,C,D,E,F$. This gives the inequality
@@
with equality if and only if $\mathrm{diam}(A)=\mathrm{diam}(B)=\dots=\mathrm{diam}(F)=0$. i.e. we have $\mathrm{len}(L^{\prime})\geq 24\pi$ with equality if and only if $L^{\prime}$ also decomposes into circles.
Subsection 7.2 is devoted to filling in the missing details. This is a very tedious endeavor included only for the sake of completeness; there are no deep insights to be gained. Therefore one might skip Subsection 7.2 to read the short proofs of the final results in Subsection 7.3.
@@ 7.2 Circle decompositions - the details
7.1 Definition (circle decomposition).
Let $C$ be a closed curve such that the definition domain of $C$ decomposes into $I_{1}\cup I_{2}\cup\dots\cup I_{n}$, where each $I_{i}$ is a compact interval of positive length with $\max(I_{i})=\min(I_{i+1})$, for all $i$ (here indices are understood to be cyclic and the definition domain is seen as an interval with endpoints identified). We define $a_{i}$ via $a_{i},a_{i+1}]=I_{i}$. Let $\{A_{j}\}_{j=1}^{m}$ and $\{\alpha_{l}\}_{l=1}^{k}$ each be a partition of $\{1,2,\dots,n\}$ into non-empty sets. We say that $\left(C,\{a_{i}\}_{i=1}^{n},\{A_{j}\}_{j=1}^{m},\{\alpha_{l}\}_{l=1}^{k}\right)$ is a circle decomposition of $C$ if for all $i_{1},i_{2}\in\{1,\dots,n\}$ and $j\in\{1,\dots,m\}$:
@@
Let $(C,\{a_{i}\},\{A_{j}\},\{\alpha_{l}\})$ be a circle decomposition. If $a_{i}$ is on the boundary of one of the intervals $I_{i^{\prime}}$ with $i^{\prime}\in A_{j}$, then we say that $a_{i}$ is along the circle $C^{j}$. We say that $\alpha_{l}$ is along the circle $C^{j}$ if there is some $i\in\alpha_{l}$ with $a_{i}$ along $C^{j}$.
@@
We say that $C$ intersects $D$ transversely at a point $p=D(x)=C(y)$ if there is an open disk $U\subseteq[0,1]^{n}$ around $x$ and an open interval $I\subseteq[0,1]$ around $y$ such that $D|_{U}$ and $C|_{I}$ only intersect in $p$ and such that the Brouwer degree of $(D-C)|_{U\times I}$ around 0 is $\pm 1$.
@@
  • $K_{j}$ circle of $C_{1}$ $\implies$ there is some $\beta_{l}$ such that $z_{j}=C_{2}(\beta_{l}),$ and every intersection of $C_{2}$ and $D_{j}$ is transverse,
  • $K_{j}$ circle of $C_{2}$ $\implies$ there is some $\alpha_{l}$ such that $z_{j}=C_{1}(\alpha_{l}),$ and every intersection of $C_{1}$ and $D_{j}$ is transverse.
@@
  • for every $A$-index $l$: $\{j\text{ }A\text{-index }\mid l\text{ along }C_{1}^{j}\}<2\#(h^{A})^{-1}(l)$,
  • for every $B$-index $l$: $\{j\text{ }B\text{-index }\mid l\text{ along }C_{2}^{j}\}<2\#(h^{B})^{-1}(l)$
@@
The main tool in proving Theorem 7.9 will be finding a lower bound for the length of almost closed curves avoiding some set of points. In the following lemmas, $U_{r}(\cdot)$ will refer to the open $r$-neighborhood of a set or point.
@@
Let $D_{1}$ and $D_{2}$ be two open disks in $\mathbb{R}^{2}$, with radii $r_{1},r_{2}$ and centers $z_{1},z_{2}$ respectively, with $r_{1},r_{2}\in[\tfrac{1}{2},1]$ and $z_{1}-z_{2}||<\tfrac{1}{2}$. For all $a,b\in\mathbb{R}^{2}\setminus\left\{D_{1}\cup D_{2}\right\}$ with $a-b||<\tfrac{1}{2}$ and all $m\in D_{1}\cap D_{2}$ not collinear with $a$ and $b$, there exists a curve $C$ from $a$ to $b$, such that
[…]
obtained by some radial projection of the straight-line segment $\overline{ab}$ from $m$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
The following theorem is an almost-closed analogon of Proposition 5.5 and an easy consequence of the previous proposition.
@@
[…]
@@
[…]
@@
Let $C_{1},C_{2},\dots,C_{n}$ be $n$ curves, all defined on compact intervals such that $p=C_{1}(x_{1})=C_{2}(x_{2})=\dots=C_{n}(x_{n})$ is the only intersection point of $C_{1}$ and $C_{2}$. Further assume that every $C_{i}$ is an embedding. For every $\varepsilon>0$ there is a $\delta>0$ such that the following holds: Let $C_{i}-\tilde{C}_{i}||<\delta$ for all $i$. Let $y_{1},\dots,y_{n}$ be such that
@@
[…]